Blue Co. makes a product with the following standard costs:

Standard Quantity or Hours

Standard Price or Rate

Direct materials

7.4 grams

$8.00 per gram

Direct labor

0.3 hours

$16.00 per hour

Variable overhead

0.3 hours

$7.00 per hour


The company reported the following results concerning this product in July.

Actual output

2,200

units

Raw materials used in production

16,420

grams

Purchases of raw materials

17,900

grams

Actual direct labor-hours

720

hours

Actual cost of raw materials purchases

$141,410

Actual direct labor cost

$12,528

Actual variable overhead cost

$5,112

The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased.

A) Compute the materials quantity variance.
B) Compute the materials price variance.
C) Compute the labor efficiency variance.

D) Compute the labor rate variance.
Answer #1

Solution :

(A) Material Quantity Variance :

Material Quantity Variance = (AQ - SQ) * SR

= (16,420 - 16,280) * $ 8

= $ 1,120 Unfavorable

(B) Material Price Variance

Material Price Variance = (SR - AR) * Actual Quantity Purchased

= ($ 8 - $ 7.90) * 17,900

= $ 1,790 Favorable

(C) Labor Efficiency Variance

Labor Efficiency Variance = (SH - AH) * SR

= (660 - 720) * $ 16

= $ 960 Unfavorable

(D) Labor Rate Variance

Labor Rate Variance = (SR - AR) * AH

= ( $ 16 - $ 17.4) * 720

= $ 1,008 Unfavorable

Working :

Standard Quantity of Material Required = 2,200 * 7.4 = 16,280

Actual Rate of Material Purchased = $ 141,410 / 17,900 = $ 7.9

Standard Hours = 2,200 * 0.3 = 660 Labor Hours

Actual Rate per Direct Labor Hour = $ 12,528 / 720 = $ 17.4 per direct labor hour

Note : Would be glad to help you further.